OpenStudy (anonymous):

11. Solve the inequality. Show your work. |r + 3| ≥ 7

OpenStudy (jdoe0001):

\(\bf |r + 3| \ge 7\implies \begin{cases} +(r + 3) \ge 7\\ \quad \\ \bf -(r + 3) \ge 7 \end{cases}\)

OpenStudy (jdoe0001):

so you'd get 2 values for "x", one for each scenario

OpenStudy (anonymous):

ohh thank you!!!

OpenStudy (jdoe0001):

yw